# Conversational SimulMT: Efficient Simultaneous Translation with Large Language Models

### v3

@@ 4 Experiments
Trajectory Dimension De-En En-Vi
Meta-Trajectory #Chunk $10.69\pm 5.51$ $12.98\pm 8.07$
#SRC word/Chunk $1.74\pm 0.78$ $1.38\pm 0.38$
#TGT word/Chunk $1.79\pm 0.78$ $1.73\pm 0.49$
Aug-Trajectory #Chunk $2.74\pm 1.21$ $3.12\pm 1.57$
#SRC word/Chunk $7.01\pm 3.92$ $5.83\pm 2.76$
#TGT word/Chunk $7.18\pm 3.90$ $7.35\pm 3.49$
Table 2: This table presents the statistics of the conversational SimulMT data curated from the training set of both benchmarks, including mean and standard deviation of the number of chunks, number of source/target tokens per chunk in the meta- and augmented-trajectory respectively.
@@ 4.1 Datasets
(4.5M training pairs) We employ newstest2013 (3000 pairs) for validation and utilize newstest2015 (2169 pairs) as our test dataset222www.statmt.org/wmt15/. For baseline models, we follow instructions in Zhang et al. (2022) to use BPE Sennrich et al. (2016) for tokenization with 32,000 merge operations, the vocabulary is shared between German and English.
@@
(133K training pairs) TED tst2012 (1553 pairs) and tst2013 (1268 pairs) are used as validation and test set333nlp.stanford.edu/projects/nmt/. We follow Ma et al. (2020); Zhang et al. (2022) to preprocess the data for baseline models, resulting in vocabulary sizes of 17K and 7.7K for English and Vietnamese.
Note that the above tokenization is only applicable for non-LLM baseline models. For LLM-based SimulMT, we use their provided vocabularies and tokenizers.
Conversational SimulMT Dataset
We employ fastalign Dyer et al. (2013) to obtain word alignment between source and target tokens. We randomly merge $\delta_{\min}=2$ to $\delta_{\max}=10$ consecutive segments of the meta trajectories. For shift operation, we set the value of $\beta$ and $\rho_{\min}$ to 0.5, i.e. with the probability of 0.5, we shift at least 50% portions of tokens in a target segment to the next one. Table 2 presents detailed statistical information about the dataset.
4.2 Evaluation Metrics.
We use SacreBLEU [^4] BLEU+nrefs:1+case:mixed+eff:no+tok:13a
@@
 (Post, 2018), COMET [^5] https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2020) and word-level average lagging (AL) (Ma et al., 2019) to evaluate the quality and latency. To measure the computational-aware latency, we use the word wall time (WWT) metric (Wang et al., 2023b), which is the average wall time to predict a word on the same hardware.
@@ 4.3 Model Training
For LLM-based methods, we follow Wang et al. (2023b) to utilize Llama2-7b-chat Touvron et al. (2023) as the backbone for both language pairs. We conduct SFT with QLoRA Hu et al. (2022); Dettmers et al. (2023) for one epoch to reduce the computation. The QLoRA adapters are configured with $r=64$ and $\alpha=16$. We set the learning rate to 2e-4, batch size to 48, and employ 4-bit quantization. For all experiments, a single A100 GPU is used. For offline prompt baselines, we tune the model on the offline bitext corpus. The LLMs using conversational prompts are tuned on our curated conversational SimulMT data (§3.2).
@@ 4.4 Settings
We compare our proposed conversational SimulMT method against the following baselines
Offline NMT
Following (Zhang and Feng, 2022), we train vanilla Transformers Vaswani et al. (2017) of size small (48M) for En->Vi and big (300M) for De->En, and use beam size of 5 during inference.
SimulMT
We report the performance of SimulMT with (i) fixed policy Wait-$k$ Ma et al. (2019) and (ii) adaptive policy, including Wait-Info Zhang et al. (2022) and ITST Zhang and Feng (2022).
  • Wait-$k$ (Ma et al., 2019) propose to achieve simultaneous decoding with an encoder-decoder Transformer by first reading $k$ source tokens and then alternately reading and writing one token per time. During decoding, we set $k$ from 1 to 8 for De->En and 4 to 8 for En->Vi.
  • Wait-Info (Zhang et al., 2022) quantitatively measures the information of each token, ensuring that the information in the target remains below the received source information with the threshold $\mathcal{K}$. We set the value of $\mathcal{K}$ from 1 to 8 for both datasets during decoding.
  • ITST (Zhang and Feng, 2022) calculates the weight of information transferred from each source token to the current target token. Based on the total information received, it then determines whether to proceed with the translation of the target token. We evaluate with the threshold ranging from 0.1 to 0.7 for both datasets in the experiment.
@@
We compare our conversational prompt against LLM-based SimulMT with offline prompt Wang et al. (2023b). For both methods, we set $n$=[3,5,7,9,11,13] in the experiment. Both approaches utilize RALCP as the stable hypothesis selection method if using beam search decoding (beam=5). Additionally, we also report GREEDY selection method (decoding with greedy search without pruning its prediction with prefix selection strategy), which serves as a low-latency baseline. LLM-based offline NMT is also reported as the performance upper bound.
Notations
For space reason, we abbreviate experimental settings in the following fields: (i) decoding mode: offline (O) or simultaneous (S), (ii) model name: dedicated SimulMT models or LLMs (LM), (iii) prompt type: offline (OP) or conversational prompt (CP) used only for LLMs, (iv) hypothesis selection method (only for LLM-based SimulMT): RALCP or GREEDY. For example, our proposed method for simultaneous decoding is written as S:LM:CP:RALCP.
Figure 3: Translation quality and latency results on two benchmarks. For abbreviations, please refer to the Notations description (§4). Results are presented in three groups with different colors: (i) specialized MT and SimulMT methods (yellow), (ii) LLM-based offline-prompting SimulMT (blue), and (iii) LLM-based conversational SimulMT (red).
@@ 4.5 Results
Our preliminary result in Table 1 have shown that out-of-the-box LLMs struggle in zero-shot and few-shot conversational SimulMT. In our experiments, we focus on evaluating the impact of fine-tuning LLMs on curated data, with the goal of improving their performance in simultaneous translation using conversational prompts. Specifically, we aim to determine if fine-tuning can enable LLMs to handle conversational SimulMT more effectively, and to explore any trade-offs between latency reduction and translation quality.
@@
Figure 3 shows the BLEU and COMET scores, as well as the average latency on two benchmarks. Overall, LLM-base SimulMT (red and blue) outperforms specialized methods (yellow), with the gain up to 3 BLEU scores and 6 COMET scores. This suggests that leveraging prior knowledge in LLMs is advantageous for translation. Furthermore, given increased latency allowance, LLM-based SimulMT can approach the performance of offline NMT. Among LLM-based methods, given the same latency, conversational SimulMT (red) achieves comparable results to its offline-prompting counterparts (blue) in BLEU and even performs better if measured by COMET. This shows our approach does not negatively impact the translation capabilities of LLMs.
@@
Compared to offline-promting method (blue), conversational SimulMT (red) can effectively reduce translation latency (1.17 and 1.50 AL in average for De->En and En->Vi) in all latency-range. It even reaches the similar latency level of specialized SimulMT (yellow) in En->Vi translation. The use of RALCP in conversational SimulMT (S:LM:CP:RALCP) can further boost quality over greedy strategy (S:LM:CP:GREEDY), though at the cost of increased latency due to its more conservative behavior. Compared to other baseline approaches, conversational SimulMT obtains a better trade-off between latency and quality, as it fits a wider latency span with more consistent quality.
More importantly, conversational SimulMT (red) can maintain better quality at the low-latency level (AL<4) compared to specialized models (yellow), making it more superior in practical application where both high quality and low latency are required. In contrast, given the same decoding configuration, decoding with offline-prompt (blue) is unable to fit into the low-latency span, making its quality advantageous less competitive compared to specialized approaches (yellow).
@@ 5 Analysis
Figure 4: This figure presents the relationship between computational-aware latency in Word Wall Time and the quality in COMET score. For settings of simultaneous decoding, they are represented by circles with the size representing the variance in both metrics under different latency control parameters such as $n$. For offline settings, diamond marks are used. Colors and notations are consistent with the Figure 3, where our proposed approach is highlighted with bold text.
@@ 5.1 Decoding Speed
While AL is effective in algorithmically quantifying the delay between the translation and the source input, it does not account for the computational cost associated with the inference process, i.e. assuming unit decoding speed across all models. In real-world applications, the actual inference wall time is a critical factor that directly impacts user experience. A model may exhibit low latency in terms of AL but incur a high inference cost, resulting in substantial time latency. To address this, we also report decoding speed using the word wall time (WWT) metric, which measures the real inference time per word (§4.4).
Detailed result in WMT15 De->En is shown in Figure 4. Overall, specialized MT and SimulMT The combination of offline-prompting and RALCP (S:LM:OP:RALCP) is the slowest configuration, significantly lagging behind its offline setting (O:LM:OP), despite of obtaining a good translation quality. Conversely, when employing conversational prompting (S:LM:CP:RALCP), it can achieve comparable decoding speed with the offline model (O:LM:CP) while maintaining the translation quality level. Moreover, employing greedy hypothesis selection with the conversational prompt (S:LM:CP:GREEDY) achieves a comparable speed to specialized SimulMT (yellow) but with superior quality.
Figure 5: The performance of the model tested with different levels of trajectory augmentation strategy.
@@ 5.2 Effectiveness of Trajectory Augmentation
We conduct an ablation study to assess the impact of trajectory augmentation methods by supervised fine-tuning LLMs on data derived from (i) meta trajectories only, (ii) incorporating merge operations, and (iii) both merge and shift operations (§3.2). The fine-tuning hyper-parameters are kept constant, with only the training data varies. Figure 5 reports the performance of conversational SimulMT with RALCP selection strategy under different training data configurations. Interestingly, compared to the model trained solely on meta trajectories, trajectory augmentation does not only enhance translation performance but also reduces latency, with the merge operation playing a pivotal role. We speculate that without augmentation, the model may have difficulty predicting accurately under varying latency conditions. This leads to RACLP accepting less stable hypotheses, hence worsening overall latency.
5.3 Ability of Context Following
Figure 6: The model performance for translating the last source chunk w/wo context under different chunk sizes.
Performing SimulMT with conversational prompts requires the LLM’s proficiency in utilizing context information accurately. To further examine the context-following capability of the model tuned with our curated data, we assess its performance in translating the final chunk of the source text, both with and without preceding context. Specifically, we extract the inference history for each instance, dividing it into a source-target dialogue history as context and the final source chunk as the latest input. Subsequently, the fine-tuned LLM is tasked with generating the last translation chunk from the last source chunk, either with or without the translation history. The evaluation is done by computing the BLEU score on the concatenation of the last chunk to its original history. A consistent 2-point decrease in performance can be found when decoding without context for all chunk sizes. This underscores the LLM’s ability to effectively harness context information for producing more precise translations, as illustrated in Figure 6.

### v4

@@ 4 Experiments
[…]
Meta-Trajectory #Chunk $10.69\pm 5.5$ $2.98\pm 8.1 11$$1.94\pm 7.3$
#SRC word/Chunk $1.74\pm 0.8$ $1.38\pm 0.4$ $1.68\pm 0.5$
#TGT word/Chunk $1.79\pm 0.8$ $1.73\pm 0.5$ $1.53\pm 0.5$
Aug-Trajectory #Chunk $2.74\pm 1.2$ $3.12\pm 1.6$ $2.95\pm 1.4$
#SRC word/Chunk $7.01\pm 3.9$ $5.83\pm 2.8$ $7.02\pm 3.6$
#TGT word/Chunk $7.18\pm 3.9$ $7.35\pm 3.5$ $6.40\pm 3.2$
[…]
@@ 4.1 Datasets
(4.5M training pairs) We use newstest2013 (3000 pairs) for validation and newstest2015 (2169 pairs) for testing222www.statmt.org/wmt15/.
@@
(133K training pairs) We employ TED tst2012 (1553 pairs) and tst2013 (1268 pairs) as validation and test sets, respectively333nlp.stanford.edu/projects/nmt/.
MUST-C En->Zh
(di-gangi-etal-2019-must) (359k training pairs) This TED talk dataset provides 1349 pairs for validation and the tst-COMMON (2841 pairs) for testing.
Conversational SimulMT Datasets
For each dataset, we create conversational prompt versions from their training sets using the approach described in §3.2. We employ fastalign fastalign to obtain initial word alignment graphs. For trajectory augmentation, we set $\delta_{\min:\max}=(2,10)$ for merging operations. For shift operations, both $\beta$ and $\rho_{\min}$ are set to 0.5, meaning we shift at least 50% of tokens in a target segment to the next one with 0.5 probability. Table˜2 presents detailed statistics for these datasets.
4.2 Evaluation Metrics
We evaluate translation quality and latency using SacreBLEU [^4] BLEU+nrefs:1+case:mixed+eff:no+tok:{13a,zh}
@@ 4.2 Evaluation Metrics
 (matt-2018-sacrebleu), COMET [^5] https://huggingface.co/Unbabel/wmt22-cometkiwi-da (rei2020comet), and word-level average lagging (AL) (ma-etal-2019-stacl). To assess computational efficiency, we measure word wall time (WWT) (wang2023simultaneous), which represents the average time required to predict a word on identical hardware.
@@ 4.3 Model Training
For all LLM-based methods, we use Llama2-7b-chat (touvron2023llama2) as the backbone following wang2023simultaneous. We conduct QLoRA-based SFT hu2022lora; dettmers2023qlora for one epoch with $r=64$, $\alpha=16$, learning rate of 2e-4, batch size of 48, and 4-bit quantization on a single A100 GPU. Both offline and conversational prompt models are fine-tuned on identical data sources (standard offline style bitext from the aforementioned training sets), but formatted as offline prompts and conversational prompts respectively.
Figure 3: Translation quality and latency results on three benchmarks. Results are presented in three groups with different colors: (i) Encoder-Decoder Transformer baselines (orange), (ii) Offline-Prompt LLMs (blue), and (iii) Conversation-Prompt LLMs (red). Offline and Simultaneous decoding are distinguished by the first letter (O/S).
@@ 4.4 Settings
We compare our proposed conversational SimulMT against the following baselines:
Encoder-Decoder Transformers
We evaluate the performance of a series of specialized Encoder-Decoder Transformer models for both OfflineMT and SimulMT:
  • Offline NMT: Following (zhang2022itst), we train vanilla Transformer vaswani2017transformer (48M parameters for En->Vi; 300M for De->En and Zh->En) with beam size 5 for inference.
  • Wait-$k$ (ma-etal-2019-stacl): A fixed policy approach that reads $k$ source tokens before alternating read/write operations. We test with $k$ ranging from 1-8 for De->En and Zh->En, 4-8 for En->Vi.
  • ITST (zhang2022itst) An adaptive policy that measures the information transferred from source to target token and determines when to proceed with translation with a threshold (set as 0.1-0.7 for all datasets).
  • Wait-Info (zhang-etal-2022-wait) An adaptive policy using token information thresholds ($\mathcal{K}$ from 1-8 for all datasets) to coordinate the timing of translation.
@@
We compare our conversational prompt approach against the offline prompt method (wang2023simultaneous), using identical READ policies with chunk sizes nn=[3,5,7,9,11,13]. Both approaches are evaluated with RALCP hypothesis selection (beam=5). We also assess greedy decoding (beam=1, no hypothesis selection) with our conversational prompting only (as computational latency baseline), since offline prompting inherently requires hypothesis selection and cannot function with greedy search. For reference, we include results from LLM-based OfflineMT as a performance upper bound.
@@ 4.5 Results
Our preliminary study in Table˜1 showed LLMs struggle with zero/few-shot conversational SimulMT. Here we examine whether fine-tuning on our curated data enables effective conversational SimulMT, focusing on quality-latency balance.
Figure 4: Relationship between computational efficiency (Word Wall Time) and translation quality (COMET score) on WMT15 De->En. Simultaneous decoding settings are shown as circles, with circle size representing variance across different latency control parameters (e.g. $n$). Offline settings are represented by diamonds. Color coding matches Figure˜3, with our proposed approach highlighted in bold.
@@
As shown in Figure˜3, LLM-based approaches (red and blue) outperform Transformer baselines (yellow) across all language pairs by up to 3 BLEU/10 COMET points. With sufficient latency allowance, LLM-based SimulMT even surpasses offline Transformer NMT. At equivalent latency levels, our conversational prompting (red) achieves comparable BLEU scores to offline prompting (blue) while often showing better COMET scores.
@@
Our conversational SimulMT (red) reduces latency compared to offline prompting (blue), with average reductions of 1.17 and 1.50 AL across all benchmarks. For En->Vi and En->Zh, our approach achieves latency comparable to specialized SimulMT models. While RALCP (S:LLM-ConvPrompt-RALCP) generally provides better quality than greedy decoding (S:LLM-ConvPrompt-Greedy), the latter offers lower latency.
Practical Advantages
Most significantly, our conversational SimulMT (red) maintains superior translation quality at low latency levels (AL<4) compared to specialized models (yellow), making it particularly valuable for practical applications requiring both high quality and low latency. In contrast, offline prompting (blue) with identical decoding configurations struggles to operate effectively in the low-latency range, diminishing its quality advantages relative to specialized approaches (yellow). These results demonstrate that our conversational prompting approach effectively addresses the efficiency-quality trade-off in simultaneous translation with LLMs.
@@ 5.1 Decoding Speed
While Average Lagging (AL) effectively quantifies algorithmic delay between translation and source input, it doesn’t account for computational costs. In real-world applications, actual inference time critically impacts user experience: a model with low AL might still deliver poor user experience due to high computational overhead. To address this limitation, we evaluate decoding speed using Word Wall Time (WWT), which measures actual inference time per word (§4.4).
Figure 4 presents detailed WWT results for WMT15 De->En translation. Our analysis reveals that offline prompting with RALCP (S:LLM-OffPrompt-RALCP) exhibits the slowest performance, making it impractical despite good translation quality. In contrast, our conversational prompting approach with RALCP (S:LLM-ConvPrompt-RALCP) achieves computational efficiency comparable to offline LLM translation (O:LLM-ConvPrompt-Beam=5) while maintaining high translation quality.
Most notably, our conversational prompting with greedy decoding (S:LLM-ConvPrompt-Greedy) delivers the best efficiency-quality balance—achieving processing speeds comparable to specialized SimulMT models (yellow) while producing significantly better translations. These results demonstrate that our approach effectively addresses both algorithmic and computational latency concerns, making it suitable for practical deployment.
Figure 5: Effect of trajectory augmentation strategies on translation quality (BLEU) and latency (AL) for WMT15 De->En. Results compare models trained on meta-trajectories alone versus with merge and shift operations.
@@ 5.2 Effectiveness of Trajectory Augmentation
To evaluate our trajectory augmentation strategy, we conducted an ablation study comparing models trained on: (i) meta trajectories only, (ii) meta trajectories with merge operations, and (iii) meta trajectories with both merge and shift operations (§3.2). All models used identical hyperparameters, with training data as the only variable.
As shown in Figure˜5, trajectory augmentation yields notable improvements in translation quality and latency when using RALCP. The merge operation contributes most significantly to these improvements, while models trained solely on meta trajectories perform poorly across all metrics.
This suggests augmentation techniques enhance the model’s ability to generalize across different latency conditions. Without augmentation, the model struggles with varying input chunk sizes, causing RALCP to accept less reliable hypotheses and increasing latency. The augmented approach effectively prepares the model for dynamic simultaneous translation scenarios.
Figure 6: Translation quality (BLEU) on WMT15 De->En when generating the final chunk with vs. without preceding context, across different chunk sizes. The consistent gap demonstrates effective context utilization.
5.3 Ability to Leverage Contextual Information
Effective SimulMT with conversational prompting requires the model’s ability to accurately utilize contextual information. To evaluate this capability, we designed an experiment isolating the model’s performance on the final chunk of translation both with and without access to preceding context.
For each test instance, we extracted the complete inference history and separated it into: (i) the source-target dialogue history serving as context, and (ii) the final source chunk representing the latest input. We then tasked our fine-tuned LLM with translating this final chunk under two conditions: with and without access to the preceding conversation history. Performance was evaluated by computing BLEU scores on the concatenation of the generated final chunk with its original history.
As shown in Figure˜6, we observed a consistent 2-point decrease in BLEU scores when context was withheld. This performance gap demonstrates our model effectively leverages information from previous conversation turns to produce more accurate translations, confirming the fine-tuned LLM maintains translation coherence.
5.4 Generalizability Across LLM Families
(a) Impact of model iteration (Llama-2-7b-chat vs. Llama-3.1-8B-Instruct) on WMT15 De->En.
(b) Effect of model scale (Llama-3.1-8B-Instruct vs. Llama-3.2-3B-Instruct) on WMT15 De->En.
(c) Impact of target language proficiency (Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct) on MUST-C En->Zh.
Figure 7: Performance comparison of different LLM families with our conversational prompt.
In our main experiments, we used Llama-2-7b-chat following wang2023simultaneous for consistency. Now, we examine our approach’s generalizability across different LLMs, using identical training and inference parameters for fair comparison. We report only greedy simultaneous decoding and offline beam=5 results to eliminate interference with hypothesis selection.
Impact of Model Iteration
We compare Llama-2-7b-chat with the newer Llama-3.1-8B-Instruct (grattafiori2024llama3herdmodels) on WMT15 De->En to assess how model advancements affect performance. As shown in Figure˜7(a), the newer model demonstrates consistent improvements in both offline and simultaneous modes. This confirms that conversational SimulMT effectively transfers to newer LLMs, with benefits from improved instruction-following capabilities and enhanced language modeling.
Effect of Model Scale
We investigate how model size impacts performance by comparing Llama-3.1-8B-Instruct with the smaller Llama-3.2-3B-Instruct (grattafiori2024llama3herdmodels) on WMT15 De->En. Figure˜7(b) shows that while the larger model predictably outperforms its smaller counterpart, the 3B model still achieves acceptable translation quality (on par with Llama-2-7b-chat in Figure˜7(a)), suggesting our method is viable on resource-constrained devices.
Impact of Target Language Proficiency
We evaluate Llama-3.1-8B-Instruct against Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport) on MUST-C En->Zh to investigate the effect of the model’s target language capabilities. As shown in Figure˜7(c), Qwen2.5 consistently outperforms Llama-3.1 for Chinese translation by 1-2 BLEU points across all latency settings, demonstrating that target language proficiency provides additional benefits with our approach.
6 Related Works
Simultaneous Machine Translation (SimulMT)
is the task to provide real-time translation of a source sentence stream where the goal is to minimize the latency while maximizing the translation quality. A common approach is to train an MT model on prefix-to-prefix dataset to directly predict target tokens based on partial source tokens (ma-etal-2019-stacl). Alternatively, low_latency proposed the incremental decoding framework to leverage the pretrained OfflineNMT model and turn it into a SimulMT model without further training. A core component of SimulMT is a read-write policy to decide at every step whether to wait for another source token (READ) or to generate a target token (WRITE). Previous methods have explored fixed policy, which always waits for $k$ tokens before generation  (ma-etal-2019-stacl; zhang-etal-2022-wait) and adaptive policy, which trains an agent via reinforcement learning (gu-etal-2017-learning; arthur-etal-2021-learning). Re-translation (DBLP:conf/acl/ArivazhaganCMCY19) from the beginning of the source sentence at the WRITE step will incur high translation latency. Stable hypothesis detection methods such as Local Agreement, hold-$n$ (low_latency) and Share prefix SP-n (nguyen21c_interspeech) are employed to commit stable hypothesis and only regenerate a subsequence of source sentence. The goal is to reduce the latency and minimize the potential for errors resulting from incomplete source sentence (polak-etal-2022-cuni; wang2021diformer).
LLM-based NMT
Recent research has delved into the potential usage of LLMs in MT DBLP:journals/corr/abs-2302-09210; DBLP:journals/corr/abs-2304-04675; DBLP:journals/corr/abs-2309-07423, especially in handling discourse phenomena (wang-etal-2023-document-level; wu2024adapting) and linguistic nuances such as idioms (manakhimova-etal-2023-linguistically) and proverbs (wang2025proverbsrunpairsevaluating). While LLMs do exhibit some level of translation capability, prior research has identified that they still lags behind the conventional NMT models, especially for low resource languages (DBLP:journals/corr/abs-2309-07423). Additionally, the translation performance varies depending on prompting strategies (Zhang2023PromptingLL). Efforts have been made to enhance the LLMs’ MT performance by incorporating guidance from dictionary (lu2023chainofdictionary), further fine-tuning (zeng2023tim; xu2023paradigmshift) and augmenting with translation memories (mu-etal-2023-augmenting).
LLM-based SimulMT
SimulLLM (agostinelli2023simul) explore the ability to adapt an LLM finetuned on NMT task to simultaneous translation with wait-k strategy. wang2023simultaneous adopt hybrid READ/WRITE policy with wait-k and incremental decoding. TransLLaMA (koshkin2024transllama) teach LLMs to produce WAIT tokens to preserve the causal alignment between source and target tokens. At each inference round, LLMs only produce a single word or WAIT token, which is very costly due to multiple rounds of LLM calls. guo2024sillm introduce LLM into the SimulMT task as a translation agent working with a specialized SimulMT policy agent. An additional memory module stores translation history. The policy agent decides on READ/WRITE actions, while the LLM translates target segments. They face the same KV-cache reuse challenge noted by wang2023simultaneous, making the computational cost of collaborating big and small models even more significant.